# A geometric perspective on Algebraic Quantum Field Theory

### v1

@@ 3.1 Non-modular covariant nets of standard subspaces
A long standing question in AQFT concerns the necessity to assume the Bisognano–Wichmann property. In particular, given an isotonous, covariant, local, (positive energy) net of von Neumann algebras, does the Bisognano–Wichmann property hold? This is always the case if the local algebra net is generated by Wightman fields as it is proved in [BW75]. It holds in conformal theories ([Lo08, BGL93]) and in massive theories [Mu01]. In some scalar case one can give an algebraic condition ensuring the Bisognano–Wichmann property that applies also in possibly interacting theories [Mo18, DM20]. When infinite particle degeneracy is present many counterexamples to Bisognano-Wichmann property arise [Mo18, MT19]. On the other hand in these counterexamples the wedge modular group implements a covariant representation of the symmetry group. In [Yn94], Yngvason provided a translation covariant von Neumann algebra net where there is no geometric modular action of the wedge algebra. On the other hand these nets are not expected to be Lorentz covariant.
Since the modular theory of a net of von Neumann algebras in the vacuum representation is contained in the standard subspace structure of wedge algebras, it is a natural problem to provide counterexamples to the Bisognano–Wichmann property in the net of standard subspace context without modular covariance on wedge subspaces. This problem has been recently faced in [MN22] providing a family of non-local standard subspace nets with no modular action of the wedge modular group. Here we point out the main steps of the construction.
Let $G=\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})$ with Lie algebra ${\mathfrak{g}}$, let $G$ be the graded extension $G_{\tau}=G\rtimes\{1,\tau\}$, where $\tau$ is an Euler involution and let $U$ be an (anti-)unitary representation of $G_{\tau}$ on a Hilbert space $\mathcal{H}$. Assume the existence of a subgroup $H_{\tau}=H\rtimes\{e,\tau\}\subseteq G_{\tau}$, two Euler wedges $W_{1}=(h_{1},\tau_{1})\in\mathcal{G}_{E}(H_{\tau})$ and $W_{2}=(h_{2},\tau_{2})\in\mathcal{G}_{E}(G_{\tau})$. Note that $h_{1}$ is not necessarily in $\mathcal{E}({\mathfrak{g}}).$ Assume that the stabilizer $H_{W_{1}}$ of $W_{1}$ in $H$ fixes $W_{2}$, in particular $[h_{1},h_{2}]=0$. We denote the orbit $H.W_{1}$ with $\mathcal{W}_{+}(H,W_{1})$.
@@
The net ${\sf N}$ is not necessarily modular covariant as the modular groups of the wedge subspaces $U(g){\sf N}_{2}$ are in general subgroups of the larger group $G.$ The following is an equivalent condition for the modular covariance property (18) (cf. [MN22, Lem 3.1]). The net ${\sf N}$ on $\mathcal{W}_{+}(H,W_{1})$ satisfies modular covariance if and only if, for all ${g\in H},t\in{\mathbb{R}}$, the operator
@@
$[{\mathfrak{h}},h_{1}-h_{2}]\not\subseteq\ker(\mathop{{\rm ad}}\nolimits\,h_{2}),$ (23)
@@
We now indicate how to use this prescription to construct a non-modular covariant net on the two-dimensional de Sitter space. Here the key choice is to take $h\in{\mathfrak{g}}$ to be a non-symmetric Euler element in ${\mathfrak{g}}$ and check that (22) is not satisfied by testing (23). We refer to [MN22] for the general discussion and further examples as non-modular covariant nets on Minkowksi space and the representation theoretic results.
Consider the case $G=\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})$ where ${\mathfrak{g}}$ a simple non-compact Lie algebra. Consider a non-symmetric Euler element $h_{2}\in{\mathfrak{g}}$ and the associated wedge $W_{2}=(h_{2},\tau_{2})\in\mathcal{G}_{E}(G_{\tau})$. By [MN22, Lem. 2.15] there exists a $\mathop{{\mathfrak{gl}}}\nolimits_{2}$-subalgebra ${\mathfrak{b}}\subseteq{\mathfrak{g}}$ containing $h_{2}$ such that ${\mathfrak{h}}:=[{\mathfrak{b}},{\mathfrak{b}}]\cong\mathop{{\mathfrak{sl}}}\nolimits_{2}({\mathbb{R}})$ satisfies $[{\mathfrak{h}},h_{2}]\neq 0$. We consider $H:=\mathop{{\rm Inn}}\nolimits_{\mathfrak{g}}({\mathfrak{h}})\cong\mathop{{\rm SL}}\nolimits_{2}({\mathbb{R}})$ and $H_{\tau_{2}}=H\rtimes\{1,\tau_{2}\}$ (see [MN22, Lem 2.15(d)]). We assume the Euler element $h_{2}\in{\mathfrak{b}}$ to have a central component $h_{c}$, so that
@@
Let $V$ be the restriction $U|_{H_{\tau_{2}}}$. The group $H\cong\mathop{{\rm SL}}\nolimits_{2}({\mathbb{R}})$ is the double covering of $\mathop{{\rm PSL}}\nolimits_{2}({\mathbb{R}})\cong\mathcal{L}_{+}^{\uparrow}$, the Lorentz group of ${\mathbb{R}}^{1,2}$. So it acts on de Sitter space $\mathop{{\rm dS}}\nolimits^{2}$ through the covering map $\Lambda:\mathop{{\rm SL}}\nolimits_{2}({\mathbb{R}})\ni g\mapsto\Lambda(g)\in\mathop{{\rm PSL}}\nolimits_{2}({\mathbb{R}})$. We thus obtain a net ${\sf H}^{\mathop{{\rm dS}}\nolimits}$ on de Sitter spacetime as follows [MN22, Thm. 3.4]: Let
@@
where ${\sf N}_{U}$ is the BGL net defined by $U$. Then the net ${\sf H}^{\mathop{{\rm dS}}\nolimits}$ is Lorentz covariant and does not satisfy modular covariance. Indeed $h_{2}-h_{1}=h_{c}-2h_{1}$ and
@@
In particular (23) is violated and ${\sf N}$ is not modular covariant. One can also see this by observing that the Lie algebra generated by $gh_{2}g^{-1}$ with $g\in\mathop{{\rm SL}}\nolimits_{2}({\mathbb{R}})$ is $\mathop{{\mathfrak{gl}}}\nolimits_{2}({\mathbb{R}})$ so that $\Delta^{it}_{{\sf H}^{\mathop{{\rm dS}}\nolimits}(gW^{\mathop{{\rm dS}}\nolimits}_{R})}=U(\exp(2\pi gh_{2}g^{-1}t))$ can not generate a representation of $\widetilde{\mathop{{\rm SL}}}\nolimits_{2}({\mathbb{R}})$. Details in [MN22, Sect. 3.2.1 and 3.2.2]. We can further comment that counterexamples satisfying Lorentz covariance, locality and violating modular covariance are not covered by these examples and are still missing.
@@ 3.2 The Euler Element Theorem
The following theorem is an important step in understanding the key role of Euler elements in these constructions and their relation with localization properties of a net of standard subspaces or von Neumann algebras. It addresses the question, what does characterize the boosts to be such a special family of transformations that allows the Bisognano–Wichmann property to start all this discussion? Given a unitary representation of a Lie group and a standard subspace ${\tt V}$ satisfying the Bisognano–Wichmann property with respect to a generic one-parameter group in $G$, the generator of the latter in $U$ is an Euler element if a regularity property is satisfied. The following Theorem ensures that choosing Euler element is the only choice under natural conditions (cf. equation (20)).
@@
(Euler Element Theorem) [MN24, Thm. 3.1] Let $G$ be a connected finite-dimensional Lie group with Lie algebra ${\mathfrak{g}}$ and $h\in{\mathfrak{g}}$. Let $(U,\mathcal{H})$ be a unitary representation of $G$ with discrete kernel. Suppose that ${\tt V}$ is a standard subspace and $N\subseteq G$ an identity neighborhood such that the following hold:
@@
    Bisognano–Wichmann property: $U(\exp(th))=\Delta_{\tt V}^{-it/2\pi}$ for $t\in{\mathbb{R}}$, i.e., $\Delta_{\tt V}=e^{2\pi i\,\partial U(h)}$.
@@
Then $h$ is an Euler element and the conjugation $J_{\tt V}$ satisfies
@@
The theorem implies that, under the above assumptions, if ${\mathfrak{g}}$ is a compact Lie algebra, then every Euler element $h\in{\mathfrak{g}}$ is central, so that ${\tau_{h}=\mathop{{\rm id}}\nolimits_{\mathfrak{g}}}$. Therefore, a standard subspace ${\tt V}$ associated to a pair $(h,\tau)\in\mathcal{G}_{E}$ by the BGL construction can only satisfy the regularity condition in Theorem 31(b) if ${\tt V}$ is $U(G)$-invariant.
@@
It is a consequence of Theorem 31 that $\tau_{h}$ integrates to an involutive automorphism $\tau_{h}$ on the group $U(G)\cong G/\ker(U)$ that is uniquely determined by
$\tau_{h}(\exp x)=\exp(\tau_{h}(x))\quad\mbox{ for }\quad x\in{\mathfrak{g}}.$
@@
In particular the theorem ensures the extendibility of $U$ to an (anti-)unitary representation of $G_{\tau_{h}}$ on the same Hilbert space if $U$ is a faithful representation of $G$ satisfying Bisognano–Wichmann and regularity properties.
@@
The regularity property is very natural in the setting of an analysis of state localization in AQFT, see for instance [BGL02, GL95, LMR16]. Firstly note that the condition (b) is equivalent to the existence of a cyclic subspace ${\sf K}\subset{\sf H}$ such that $U(g){\sf K}\subset{\tt V}$ for every $g\in N$. This corresponds to a localization property of standard subspace nets that are finer than wedges. To provide examples we firstly refer to theories on Minkowski spacetime. We call an open subset $\mathcal{O}\subset{\mathbb{R}}^{1,d}$ spacelike if $x_{0}^{2}<{\bf{x}}^{2}$ holds for all $(x_{0},{\bf{x}})\in\mathcal{O}$ and a spacelike open subset is called a spacelike (convex) cone if, in addition, it is a pointed convex cone. Given a spacelike cone $\mathcal{C}$ with apex in $a\in{\mathbb{R}}^{1,d}$ then $\mathcal{C}=\bigcap\{gW_{R}+a:g\in\mathcal{L}_{+}^{\uparrow},gW_{R}\supset\mathcal{C}-a\}$ and for every wedge $W$ there always exists a spacelike cone $\mathcal{C}\subset gW$, for every $g$ in some small enough neighbourhood of the identity $N\subset\mathcal{P}_{+}^{\uparrow}$. Let ${\sf H}$ be a net of real subspaces on open regions satisfying isotony, covariance, cyclicity on wedge regions, then the assumption (a) and (b) in the theorem correspond respectively to the Bisognano–Wichmann property and the cyclicity for ${\sf H}(\mathcal{C})$. This is always the case for the free one-particle nets on Minkowksi space when $\mathcal{C}$ is a spacelike cone, see [BGL02, Sect. 4]. One can make an analogue remark for de Sitter spacetime. One can consider the intersection of pointed convex spacelike cones with apex in the origin in ${\mathbb{R}}^{1,d}$ with de Sitter space $\rm{dS}^{d}$ and $\mathcal{C}\cap\rm{dS}^{d}$ are bounded regions obtained by wedge intersection.
@@
Another remark, in view of [DM20] there exists no net of standard subspaces on spacelike cone regions undergoing an irreducible massless finite (non-zero) helicity $\mathcal{P}_{+}^{\uparrow}$-representation $U$. The argument can be seen as a consequence of Theorem 31 since $U$ should then extend to a representation of $\mathcal{P}_{+}$ on the same Hilbert space. This is not possibile for finite helicity representation, cf. [Va85, Thm. 9.10].
@@
The Euler Element Theorem can be applied in the setting of nets of standard subspaces on causal manifolds, see [MN24, Thm. 3.4]. With notations as in Sect. 2.4.3, let $(U,\mathcal{H})$ be a unitary representation of the connected Lie group $G$ with $\ker(U)$ discrete. If $({\sf H}(\mathcal{O}))_{\mathcal{O}\subseteq M}$ is a net of real subspaces on (the open subsets of) a $G$-manifold $M$ that satisfies (Iso), (Cov), (RS) and (BW), then Theorem 31 applies to ${\sf H}(W)$, and the Lie algebra element $h$ is an Euler element, and the conjugation $J:=J_{{\sf H}(W)}$ satisfies (27).
@@
This is a conclusion obtained by applying Theorem 31 to the standard subspace ${\sf H}=\overline{\mathcal{M}_{sa}\Omega}$, see [MN24, Thm. 3.7].
Given a local, Lorentz covariant, weakly additive net of von Neumann algebras on bounded regions of de Sitter space the authors of [BB99, Theorem 6.2], starting from the assumption that vacuum states in de Sitter space, look for any geodesic observer like equilibrium states with some a priori arbitrary temperature, show that the geodesic temperature of vacuum states has to be equal to the Gibbons–Hawking temperature. This corresponds to the fact that, under the Bisognano–Wichmann and regularity properties, the boost generator is an Euler element and its associated one-parameter group - properly parameterized - implements the modular group through the representation $U$. In the positive energy framework an analogous statement has been proved for the vacuum state in [St08, Thm. 1]. The Euler Element Theorem gives a standard subspace reformulation of these results [MN24, Cor. 3.3 and 3.5].

### v2

@@ 3.1 Non-modular covariant nets of standard subspaces
A long standing question in AQFT concerns the necessity to assume the Bisognano–Wichmann property. In particular, given an isotonous, covariant, local, (positive energy) net of von Neumann algebras, does the Bisognano–Wichmann property hold? This is always the case if the local algebra net is generated by Wightman fields as it is proved in [BW75]. It holds in conformal theories ([Lo08, BGL93]) and in massive theories [Mu01]. One can give an algebraic sufficient condition on the covariant representation ensuring the Bisognano–Wichmann property for one-particle nets that can be applied to prove the Bisognano-Wichmann property in some possibly interacting theories [Mo18, DM20]. When infinite particle degeneracy is present many counterexamples to Bisognano-Wichmann property arise [Mo18, MT19]. On the other hand in these counterexamples the wedge modular groups implement a covariant representation of the symmetry group. In [Yn94], Yngvason provided a translation covariant von Neumann algebra net where there is no geometric modular action of the wedge algebra. On the other hand these nets are not expected to be Lorentz covariant. Further developments about the relation between Modular Action, Wedge Duality and Lorentz Covariance for generalized free fields are in [GY00].
Since the modular theory of a net of von Neumann algebras in the vacuum representation is encoded in the standard subspace structure of wedge algebras, it is a natural problem to provide counterexamples to the Bisognano–Wichmann property in the net of standard subspaces context without the modular covariance property of wedge subspaces. This problem has been recently faced in [MN22] providing a family of non-local standard subspace nets with no geometric action of the wedge modular groups. We summarize below the main steps of the construction.
Let $G=\mathop{{\rm Inn}}\nolimits({\mathfrak{g}})$ where ${\mathfrak{g}}$ is simple, let $G_{\tau}=G\rtimes\{1,\tau\}$, where $\tau$ is an Euler involution, and let $U$ be an (anti-)unitary representation of $G_{\tau}$ on a Hilbert space $\mathcal{H}$. Assume the existence of a subgroup $H_{\tau}=H\rtimes\{e,\tau\}\subseteq G_{\tau}$, two Euler wedges $W_{1}=(h_{1},\tau_{1})\in\mathcal{G}_{E}(H_{\tau})$ and $W_{2}=(h_{2},\tau_{2})\in\mathcal{G}_{E}(G_{\tau})$. Note that $h_{1}$ is not necessarily in $\mathcal{E}({\mathfrak{g}}).$ Assume that the stabilizer $H_{W_{1}}$ of $W_{1}$ in $H$ fixes $W_{2}$, in particular $[h_{1},h_{2}]=0$. We denote the orbit $H.W_{1}$ with $\mathcal{W}_{+}(H,W_{1})$.
@@
The net ${\sf N}$ is not necessarily modular covariant. The following is an equivalent condition for the modular covariance property (21) (cf. [MN22, Lem 3.1]). The net ${\sf N}$ on $\mathcal{W}_{+}(H,W_{1})$ satisfies modular covariance if and only if, for all ${g\in H},t\in{\mathbb{R}}$, the operator
@@
$[h_{1}-h_{2},{\mathfrak{h}}]\not\subseteq\ker(\mathop{{\rm ad}}\nolimits\,h_{2}),$ (26)
@@
We now indicate how to use this prescription to construct a non-modular covariant net on the two-dimensional de Sitter space. The crucial step here is to choose $h\in\mathfrak{g}$ as a non-symmetric Euler element in $\mathfrak{g}$, and to verify the failure of condition (25) by testing condition (26). We refer to [MN22] for the general discussion and further examples as non-modular covariant nets on Minkowksi space and the representation theoretic results.
In the context described above, we fix ${\mathfrak{g}}$ to be a simple non-compact real Lie algebra. Consider a non-symmetric Euler element $h_{2}\in{\mathfrak{g}}$ and the associated wedge $W_{2}=(h_{2},\tau_{2})\in\mathcal{G}_{E}(G_{\tau_{2}})$. By [MN22, Lem. 2.15] there exists a $\mathop{{\mathfrak{gl}}}\nolimits_{2}$-subalgebra ${\mathfrak{b}}\subseteq{\mathfrak{g}}$ containing $h_{2}$ such that ${\mathfrak{h}}:=[{\mathfrak{b}},{\mathfrak{b}}]\cong\mathop{{\mathfrak{sl}}}\nolimits_{2}({\mathbb{R}})$ satisfies $[{\mathfrak{h}},h_{2}]\neq 0$. We consider $H:=\mathop{{\rm Inn}}\nolimits_{\mathfrak{g}}({\mathfrak{h}})\cong\mathop{{\rm SL}}\nolimits_{2}({\mathbb{R}})$ and $H_{\tau_{2}}=H\rtimes\{1,\tau_{2}\}$ (see [MN22, Lem 2.15(d)]). The Euler element $h_{2}\in{\mathfrak{b}}$ has a central component $h_{c}$, so that
@@
Let $V$ be the restriction $U|_{H_{\tau_{2}}}$. The group $H\cong\mathop{{\rm SL}}\nolimits_{2}({\mathbb{R}})$ is the double covering of $\mathop{{\rm PSL}}\nolimits_{2}({\mathbb{R}})\cong\mathcal{L}_{+}^{\uparrow}$, the Lorentz group of ${\mathbb{R}}^{1,2}$. So it acts on de Sitter space $\mathop{{\rm dS}}\nolimits^{2}$ through the covering map $\Lambda:\mathop{{\rm SL}}\nolimits_{2}({\mathbb{R}})\ni g\mapsto\Lambda(g)\in\mathop{{\rm PSL}}\nolimits_{2}({\mathbb{R}})$. We thus obtain a net ${\sf H}^{\mathop{{\rm dS}}\nolimits}$ on de Sitter spacetime as follows: Let
@@
where ${\sf N}_{U}$ is the BGL net defined by $U$ [MN22, Thm. 3.4]. Then the net ${\sf H}^{\mathop{{\rm dS}}\nolimits}$ is Lorentz covariant and does not satisfy modular covariance. Indeed $h_{2}-h_{1}=h_{c}-2h_{1}$ and
@@
In particular (26) is violated since $\mathop{{\rm ad}}\nolimits h_{2}|_{{\mathfrak{h}}}=-\mathop{{\rm ad}}\nolimits h_{1}$ and ${\sf N}$ is not modular covariant. One can also see this by observing that the Lie algebra generated by $gh_{2}g^{-1}$ with $g\in\mathop{{\rm SL}}\nolimits_{2}({\mathbb{R}})$ is $\mathop{{\mathfrak{gl}}}\nolimits_{2}({\mathbb{R}})$ so that $\Delta^{it}_{{\sf H}^{\mathop{{\rm dS}}\nolimits}(gW^{\mathop{{\rm dS}}\nolimits}_{R})}=U(\exp(-2\pi gh_{2}g^{-1}t))$ can not generate a representation of $\widetilde{\mathop{{\rm SL}}}\nolimits_{2}({\mathbb{R}})$. Details in [MN22, Sect. 3.2.1 and 3.2.2]. We can further comment that counterexamples satisfying Lorentz covariance, locality and violating modular covariance are not covered by these examples and are still missing.
@@ 3.2 The Euler Element Theorem
The following theorem is an important step in the understanding the key role of Euler elements in these constructions. It addresses the question: Are Euler elements necessary for defining fundamental localization regions in one-particle and von Neumann algebra nets? Given a unitary representation of a Lie group and a standard subspace ${\tt V}$ satisfying the Bisognano–Wichmann property with respect to a generic one-parameter group in $G$, the generator of the latter in $U$ is an Euler element if a regularity property is satisfied.
@@
(Euler Element Theorem) [MN24, Thm. 3.1] Let $G$ be a connected finite-dimensional Lie group with Lie algebra ${\mathfrak{g}}$ and $h\in{\mathfrak{g}}$. Let $(U,\mathcal{H})$ be a unitary representation of $G$ for which
$\ker({\tt d}U)\cap[h,{\mathfrak{g}}]=\{0\}.$
Suppose that ${\tt V}$ is a standard subspace and $N\subseteq G$ an identity neighborhood such that the following hold:
@@
    Bisognano–Wichmann property: $U(\exp(th))=\Delta_{\tt V}^{-it/2\pi}$ for $t\in{\mathbb{R}}$.
@@
Then $h$ is an Euler element (or central) and the conjugation $J_{\tt V}$ satisfies
@@
The theorem implies that, under the above assumptions, if ${\mathfrak{g}}$ is a compact Lie algebra, then $h\in{\mathfrak{g}}$ is central, so that ${\tau_{h}=\mathop{{\rm id}}\nolimits_{\mathfrak{g}}}$. Therefore, in this case, a standard subspace ${\tt V}$ associated to a pair $(h,\tau)\in\mathcal{G}_{E}$ by the BGL construction can only satisfy the regularity condition in Theorem 31(b) if ${\tt V}$ is $U(G)$-invariant.
@@
It is a consequence of Theorem 31 that $\tau_{h}$ integrates to an involutive automorphism $\tau_{h}^{U(G)}$ on the group $U(G)\cong G/\ker(U)$ that is uniquely determined by
$\tau_{h}^{U(G)}(\exp x)=\exp(\tau_{h}(x))\quad\mbox{ for }\quad x\in{\mathfrak{g}}.$
@@
In particular the theorem ensures the extendibility of $U$ to an (anti-)unitary representation of $G_{\tau_{h}^{G}}$ on the same Hilbert space if $U$ is a faithful representation of $G$ satisfying Bisognano–Wichmann and regularity properties.
@@
The regularity property is very natural in the setting of an analysis of state localization in AQFT, see for instance [BGL02, GL95, LMR16]. Firstly note that the condition (b) is equivalent to the existence of a cyclic subspace ${\sf K}\subset{\sf H}$ such that $U(g){\sf K}\subset{\tt V}$ for every $g\in N$. This corresponds to a localization property of standard subspace nets that are finer than wedges. To provide examples we firstly refer to theories on Minkowski spacetime. We call an open subset $\mathcal{O}\subset{\mathbb{R}}^{1,d}$ spacelike if $x_{0}^{2}<{\bf{x}}^{2}$ holds for all $(x_{0},{\bf{x}})\in\mathcal{O}$. Here we will consider spacelike cones that are also pointed, convex and causally complete cones, namely $\mathcal{C}=\mathcal{C}^{\prime\prime}$. Given a spacelike cone $\mathcal{C}$ with apex in $a\in{\mathbb{R}}^{1,d}$ then $\mathcal{C}=\bigcap\{gW_{R}+a:g\in\mathcal{L}_{+}^{\uparrow},gW_{R}\supset\mathcal{C}-a\}$ and for every wedge $W$ there always exists a spacelike cone $\mathcal{C}\subset gW$, for every $g$ in some small enough neighbourhood of the identity $N\subset\mathcal{P}_{+}^{\uparrow}$. Let ${\sf H}$ be a net of real subspaces on open regions satisfying isotony, covariance, cyclicity on wedge regions, then the assumption (a) and (b) in the theorem correspond respectively to the Bisognano–Wichmann property and the cyclicity for ${\sf H}(\mathcal{C})$. The latter property is satisfied by the free one-particle nets on Minkowksi space when $\mathcal{C}$ is a spacelike cone, see [BGL02, Sect. 4]. One can make an analogue remark for de Sitter spacetime. One can consider the intersection of pointed convex spacelike cones with apex in the origin in ${\mathbb{R}}^{1,d}$ with de Sitter space $\rm{dS}^{d}$ and $\mathcal{C}\cap\rm{dS}^{d}$ are bounded regions obtained by wedge intersection.
@@
Another remark, in view of [DM20] there exists no net of standard subspaces on spacelike cone regions, satisfying the Bisognano-Wichmann property for wedge subspaces, undergoing an irreducible massless finite (non-zero) helicity $\mathcal{P}_{+}^{\uparrow}$-representation $U$. By contradiction, if there exists such a net, then Theorem 31 applies and $U$ should extend to a representation of $\mathcal{P}_{+}$ on the same Hilbert space. This is never the case for finite non-zero helicity representation, cf. [Va85, Thm. 9.10].
@@
The Euler Element Theorem can be applied to nets of standard subspaces on causal manifolds, see [MN24, Thm. 3.4]. With notations as in Sect. 2.4.3, let $(U,\mathcal{H})$ be a unitary representation of the connected Lie group $G$ with $\ker(U)$ discrete. If $({\sf H}(\mathcal{O}))_{\mathcal{O}\subseteq M}$ is a net of real subspaces on (the open subsets of) a $G$-manifold $M$ that satisfies (Iso), (Cov), (RS) and (BW), then Theorem 31 applies to ${\sf H}(W)$, the Lie algebra element $h$ is an Euler element, and the conjugation $J:=J_{{\sf H}(W)}$ satisfies (30).
@@
This follows by applying Theorem 31 to the standard subspace ${\sf H}=\overline{\mathcal{M}_{sa}\Omega}$, see [MN24, Thm. 3.7].
Borchers and Buchholz in [BB99, Theorem 6.2] proved that given a local, Lorentz covariant, weakly additive net of von Neumann algebras on bounded regions of de Sitter space, assuming that vacuum state looks for any observer in a wedge moving through the corresponding boost-time evolution like equilibrium state with some apriori arbitrary temperature, namely a KMS-state, then the geodesic temperature of vacuum state has to be equal to the Gibbons–Hawking temperature. This corresponds to the fact that the boost generator - properly parametrized - is an Euler element. A first geometric approach to the [BB99] result is in [St08]. The Euler Element Theorem gives a standard subspace reformulation of this picture [MN24, Cor. 5.16, 3.3 and 3.5].